# arXiv:2603.03264v1[astro-ph.SR] 03 Mar 2026

### 03 Mar 2026 v1

@@ 4.1 Application to simulated spectra
We construct series of ‘synthetic observations’, relying on synthetic spectra computed with ZeeTurbo. We rely on the results of Cristofari et al. (2025a) to extract realistic filling factor distributions for series of observation nights. Consequently, we built synthetic observations from linear combinations of models computed for magnetic field strengths ranging from 0 to 10 kG in steps of 2 kG. Additionally, we rely on temperature variation estimates ($\delta T$) obtained with the LBL package (Artigau et al., 2024) from the same data to simulate the impact of temperature variations at the stellar surface, shown to be significantly anti-correlated with surface averaged magnetic field variations ($\delta\langle B\rangle$) for the stars considered in this work (Cristofari et al., 2025a). The modeled spectra are convolved with a Gaussian profile of full width at half maximum (FWHM) of 4.3 $\rm km\,s^{-1}$ to mimic the instrumental width of SPIRou, and a second Gaussian profile of 5.0 $\rm km\,s^{-1}$ to represent the typical broadening associated with rotation and macroturbulence for a star like EV Lac (see Sec. 5). We note that convolutions only approximately reproduce the impact of rotation or macro-turbulence on spectra, especially for lines with complex Zeeman patterns. The current method, however, relies on the relative variation of fluxes as a function of field strength, mitigating the impact of the approximation. Relying on convolution allows for faster computations, and our method proves relatively insensitive to broadening (see Sec. 4.4). Detailed disk integration could be added in future implementations. The broadened synthetic spectra are then re-binned on a wavelength grid typical for spectra recorded with SPIRou or Narval. Additional series of models were obtained by considering either magnetic field variations or temperature variations alone. Our process was applied to the synthetic observations, relying on different assumptions to assess the behavior of our method and its sensitivity to various parameters.
@@ 4.2 Sensitivity to normalization
Figure 1: Comparison between the retrieved (y-axis) and input (x-axis) relative magnetic field strengths ($\delta\langle B\rangle$). The synthetic observations were normalized to the template using a rolling median with a window of 100 $\rm km\,s^{-1}$. The red squares and green plus symbols (‘+’) show the results obtained before and after re-normalizing the models used to computed $\@vec{F}_{k}-\@vec{F}_{0}$ (equation 4) with a 100 $\rm km\,s^{-1}$ median filter, respectively (see text). The black line marks the equality. This figure illustrates the need for careful re-normalization of the models to avoid biases in the results.
Low-frequency fluctuations are often observed in echelle spectra, with the potential to lead to biased $\delta\langle B\rangle$ and $\delta T$ estimates. While absolute normalization of the continuum is not necessary, relative normalization with respect to the reference spectrum is crucial to obtain accurate results with the proposed method.
To mitigate the impact of low-frequency continuum fluctuations, a moving median with a width of 500 $\rm km\,s^{-1}$ was applied to the residuals between each observed spectrum and the template. The resulting continuum variations are used to correct each individual observation. We tested the impact of such a normalization step by applying it to our series of synthetic observations. Specifically, we applied rolling medians with different window sizes to synthetic observations, and applied our process with and without filtering the models used to compute $\@vec{F}_{k}-\@vec{F}_{0}$. Our simulations reveal that if our process is applied to median-filtered observations without filtering the models, the amplitude of the retrieved fluctuations are ”damped”: the structure of the variations is conserved, but their amplitude is reduced (see Fig. 1, red squares). This effect can be mitigated by filtering the models used to compute $\@vec{F}_{k}-\@vec{F}_{0}$ with a rolling median of similar window size (see Fig. 1, green ‘+’ symbols). We note that such filtering is not perfect, in particular when applied to spectral regions containing broad or blended lines. We found that applying a rolling median to the residuals between the models and model of reference (e.g. to $\@vec{F}_{k}-\@vec{F}_{0}$ directly) provides the most robust results, as this limits the changes in pseudo-continuum caused by broad absorption features. Furthermore, we note that strong magnetic fields can significantly broaden the absorption features, fundamentally changing the shape of observed pseudo-continuum. Consequently, one may favor larger windows for the rolling median in order to avoid introducing spurious feature variations in the models.
@@ 4.3 Relation between temperature and magnetic fields
Recent results provided strong evidence of anti-correlations between the obtained small-scale field measurements and temperature variations ($\delta T$) for a few strongly magnetic targets (Artigau et al., 2024; Cristofari et al., 2025a). With this work, we investigate the impact of temperature variations on magnetic field estimates, and vice-versa.
To retrieve temperature variations, we implement a process similar to that described by Artigau et al. (2024), with the notable difference that we derive $\partial\@vec{A}/\partial T$ for each pixel from our synthetic spectra. To validate our process, we created a series of modeled observations with varying effective temperatures, following the estimate obtained with the LBL approach for EV Lac (with typical variations of less than 20 K). Our implementation allows us to successfully retrieve the temperature variations used to generate the models.
To investigate the impact of magnetic fields on temperature variation estimates, we modeled spectra accounting for both temperature and magnetic fields variations taking realistic estimates for EV Lac from Cristofari et al. (2025a). The average surface magnetic field strengths then range from 4 to 5 kG, and the effective temperatures range from 3290 to 3310 K. We find that temperature variations have limited impact on the derived $\delta\langle B\rangle$. The inclusion of magnetic fields, however, has a much more significant impact on the derived $\delta T$ (see Fig. 2). We find that this impact is particularly significant for simulations covering the Narval/ESPaDOnS wavelength domain, where a very large number of atomic features are predicted to be sensitive to magnetic fields.
The results of this experiment show that $\delta T$ and $\delta\langle B\rangle$ estimates are not completely independent, with a larger sensitivity of $\delta T$ estimates to magnetic field variations for realistic conditions. We note that typical $\delta T$ are small, with a few kelvin variations leading to small changes in the observed spectra, while $\delta\langle B\rangle$ fluctuations in stars like EV Lac can lead to more visible pixel variations. Furthermore, magnetic field variations in stellar spectra lead to effects that are strongly dependent on the considered spectral line. Having both magnetically sensitive and insensitive lines in the spectral domain allows one to obtain robust constraints on $\delta\langle B\rangle$. A comparison between the model-driven and data-driven $\delta T$ estimates is discussed in Section 5.3.
@@ 4.4 Impact of atmospheric parameters on $\langle B\rangle$ variations
We investigated the impact that incorrect assumptions on the atmospheric properties of the star can have on the results. To achieve this, modeled observations were analyzed with $\@vec{F}_{k}-\@vec{F}_{0}$ computed for different effective temperatures ($T_{\rm eff}$), surface gravities ($\log{g}$), metallicities ($\rm[M/H]$) and projected rotational velocity ($v\sin{i}$).
We find that an offset in atmospheric parameters has a ‘damping’ effect on the retrieved relative field strengths, similar to the impact of short-window normalization (see Sec 4.2): the structure of the variations are typically conserved, while the amplitude of the field variations are either lower or higher than expected. Through simulations, we find that a change of $300$ K can yield a maximum difference in $\delta\langle B\rangle$ of 0.1 kG in the case of EV Lac (whose magnetic field is able to change by up to 1 kG in our simulations), comparable to the typical uncertainties reported in Cristofari et al. (2025a). Similarly, changes of 0.5 dex in the assumed surface gravity or metallicity can lead to maximum differences in $\delta\langle B\rangle$ of $0.2$ kG. Changes in the assumed broadening of the spectra of up to 5 $\rm km\,s^{-1}$ typically lead to maximum changes in $\delta\langle B\rangle$ of $0.05$ kG. The typical uncertainties on $T_{\rm eff}$, $\log{g}$, $\rm[M/H]$ and $v\sin{i}$ are generally lower than 300 K, 0.5 dex or 5 $\rm km\,s^{-1}$, and the maximum differences in $\delta\langle B\rangle$ are then lower than the typical uncertainties associated with the measurements. Our process remains therefore robust to changes in atmospheric parameters and broadening, limiting the impact of incorrect stellar characterization on the results.
@@ 5.1 Application to SPIRou spectra
We apply our process to spectra recorded with SPIRou for EV Lac, DS Leo and Barnard’s star, whose small-scale magnetic fields were previously characterized from fits of synthetic spectra to each individual observation (Cristofari et al., 2025a). This allows us to carry out direct comparisons between the obtained $\delta\langle B\rangle$ and the magnetic field estimates reported by Cristofari et al. (2025a). For each star, we assume atmospheric parameters close to those obtained by Cristofari et al. (2025a), which we list in Table 1. The broadening parameters (turbulence, rotation, instrumental width) are approximated by a single Gaussian profile. While fine tuning of each of these parameters is possible, we choose to keep the assumed atmospheric parameters and broadening kernels approximate, given the relative insensitivity of the proposed method to detailed stellar characterization (see Sec. 4.4).
Table 1: Atmospheric parameters adopted in the current work. [^1] The second column show the parameters reported by Cristofari et al. (2025a). The third column presents the approximate parameters adopted for our work, corresponding the closest point in our grid of models.
@@
For all three stars, we obtained magnetic field variations in very good agreement with the results of Cristofari et al. (2025a) (see Figs. 3 and 8), with reduced $\chi^{2}$ between the two series of points of 1.01, 1.07 and 1.26, and with Pearson correlation coefficients of $0.97$, $0.89$ and $0.58$ for EV Lac, DS Leo and Barnard’s star, respectively. Relying on the same tools as Cristofari et al. (2025a), we fit a Quasi-Periodic Gaussian Process to the retrieved $\delta\langle B\rangle$ for each star. Specifically, we rely on a Gaussian Process (GP) regression with a kernel used in previous studies (Angus et al., 2018; Fouqué et al., 2023; Cristofari et al., 2025a):
$\kappa(t_{i},t_{j})=\alpha^{2}\exp\Big[-\frac{(t_{i}-t_{j})^{2}}{2l^{2}}-\frac{1}{2\beta^{2}}\sin^{2}\Big(\frac{\pi(t_{i}-t_{j})}{P_{\rm rot}}\Big)\Big]+\sigma^{2}\delta_{ij}$ (5)
where $\alpha$ is the amplitude of the GP, $\beta$ a smoothing factor, $l$ the decay time, $t_{i}-t_{j}$ is the time difference between two data points i and j, and $\sigma$ the standard deviation of an added uncorrelated white noise. $P_{\rm rot}$ is the recurrence period, which we identify as the rotation period of the star.
The posterior distributions provide constraints on the rotation period of the star in excellent agreement with the results of Cristofari et al. (2025a), and similar long-term fluctuations are observed. These results demonstrate our ability to estimate $\langle B\rangle$ variations which are consistent with those obtained from fit of synthetic spectra to each individual observation (see Table 7 and Figs. 12, 13 and 14). Because the proposed method to derive $\delta\langle B\rangle$ relies on solving the linear equation 4 and does not require to explore the parameter space with a Markov chain Monte Carlo (MCMC), these results are obtained orders of magnitudes faster than those derived by Cristofari et al. (2025a).
Figure 3: Comparison between $\delta\langle B\rangle$ obtained in this work and the values of Cristofari et al. (2025a). The average $\langle B\rangle$ was removed from the estimates of Cristofari et al. (2025a) for comparison. The red line shows the equality. The Pearson correlation coefficient ($\rho$) is shown on the figure. Our process allows us to retrieve $\delta\langle B\rangle$ in very good agreement with those obtained from fits of synthetic spectra to the data.
@@ 5.2 Application to ESPaDOnS and Narval spectra
To complete our analysis, we apply our process to optical spectra recorded with Narval and ESPaDOnS for EV Lac and DS Leo (see Sec. 2). The derived relative magnetic field strengths are well modeled by a Quasi-Periodic Gaussian Process (see Figs. 4 and 5), yielding rotation periods of $P_{\rm rot}=4.367\pm 0.023$ d and $P_{\rm rot}=14.105\pm 0.406$ d for EV Lac and DS Leo, respectively. These estimates are in very good agreement with those reported by Cristofari et al. (2025a) of $P_{\rm rot}=4.362\pm 0.001$ d and $P_{\rm rot}=13.953\pm 0.088$ d, although with larger error bars. The larger error bars can partly be attributed to the sparsity of the available Narval and ESPaDOnS data, as a result of the observing strategies primarily focused on Zeeman Doppler Imaging reconstructions (see, e.g., Morin et al., 2008, 2011; Donati et al., 2008; Bellotti et al., 2024). This underscores the benefits of long-term spectroscopic monitoring of M dwarfs, not only for planet search, but for the study of their magnetic fields as well. We note that DS Leo was reported to exhibit differential rotation, with distinct rotation periods at the equator and the poles ($13.37\pm 0.86$ d and $14.96\pm 1.25$ d, respectively; Hébrard et al., 2016), which could in part account for the larger error bar derived with our Gaussian Process fit.
Figure 4: Best Quasi Periodic Gaussian Process (GP) fit (green) obtained on relative small-scale magnetic field estimates derived from ESPaDOnS and Narval spectra for EV Lac. The top panel shows the fit over the entire data set, while the middle and bottom panels show segments of the data set. The residuals (Res.) between data points and GP fit are shown on each panel.
Figure 5: Same as Fig. 4 for DS Leo.

### 26 Mar 2026 v2

@@ 4.1 Application to simulated spectra
We constructed a series of ”synthetic observations,” relying on synthetic spectra computed with ZeeTurbo. We relied on the results from Cristofari et al. (2025) to extract realistic filling-factor distributions for the relevant series of observation nights. Consequently, we built synthetic observations from linear combinations of models computed for magnetic field strengths ranging from 0 to 10 kG in steps of 2 kG. Additionally, we relied on temperature variation estimates ($\delta T$) obtained with the LBL package (Artigau et al., 2024) from the same data to simulate the impact of temperature variations at the stellar surface, which have been shown to be significantly anti-correlated with surface averaged magnetic field variations ($\delta\langle B\rangle$) for the stars considered in this work (Cristofari et al., 2025). The modeled spectra were convolved with a Gaussian profile of full width at half maximum (FWHM) of 4.3 $\rm km\,s^{-1}$ (to mimic the instrumental width of SPIRou), followed by a second Gaussian profile of 5.0 $\rm km\,s^{-1}$ to represent the typical broadening associated with rotation and macroturbulence for a star such as EV Lac (see Sect. 5). We note that convolutions can only approximately reproduce the impact of rotation or macro-turbulence on spectra, especially for lines with complex Zeeman patterns. The current method, however, relies on the relative variation of fluxes as a function of field strength, mitigating the impact of the approximation. Relying on convolution enables faster computations and our method proves relatively insensitive to broadening (see Sect. 4.4). We note that detailed disk integration could be added in future implementations. The broadened synthetic spectra were then re-binned on a wavelength grid typical for spectra recorded with SPIRou or Narval. Additional series of models were obtained by considering either magnetic field variations or temperature variations alone. Our process was applied to the synthetic observations, relying on different assumptions to assess the behavior of our method and its sensitivity to various parameters.
@@ 4.2 Sensitivity to normalization
Figure 1: Comparison between the retrieved (y-axis) and input (x-axis) relative magnetic field strengths ($\delta\langle B\rangle$). The synthetic observations were normalized to the template using a rolling median with a window of 100 $\rm km\,s^{-1}$. The red squares and green plus symbols (‘+’) show the results obtained before and after re-normalizing the models used to computed $\@vec{F}_{k}-\@vec{F}_{0}$ (Eq. 4) with a 100 $\rm km\,s^{-1}$ median filter, respectively (see text). The black line marks the equality. This figure illustrates the need for careful re-normalization of the models to avoid biases in the results.
Low-frequency fluctuations are often observed in echelle spectra, with the potential to lead to biased $\delta\langle B\rangle$ and $\delta T$ estimates. While an absolute normalization of the continuum is not necessary, relative normalization with respect to the reference spectrum is crucial to obtain accurate results with the proposed method.
To mitigate the impact of low-frequency continuum fluctuations, a moving median with a width of 500 $\rm km\,s^{-1}$ was applied to the residuals between each observed spectrum and the template. The resulting continuum variations were used to correct each individual observation. We tested the impact of such a normalization step by applying it to our series of synthetic observations. Specifically, we applied rolling medians with different window sizes to synthetic observations and applied our method with and without filtering the models used to compute $\@vec{F}_{k}-\@vec{F}_{0}$. Our simulations reveal that if our process is applied to median-filtered observations without filtering the models, the amplitude of the retrieved fluctuations are ”damped”, meaning that the structure of the variations is conserved, but their amplitude is reduced (see Fig. 1, red squares). This effect can be mitigated by filtering the models used to compute $\@vec{F}_{k}-\@vec{F}_{0}$ with a rolling median of similar window size (see Fig. 1, green + symbols). We note that such filtering is not perfect, in particular when applied to spectral regions containing broad or blended lines. We found that applying a rolling median to the residuals between the models and model of reference (e.g., to $\@vec{F}_{k}-\@vec{F}_{0}$ directly) provides the most robust results, as this limits the changes in pseudo-continuum caused by broad absorption features. Furthermore, we note that strong magnetic fields can significantly broaden the absorption features, fundamentally changing the shape of observed pseudo-continuum. Consequently, one may favor larger windows for the rolling median to avoid introducing spurious feature variations in the models.
@@ 4.3 Relation between temperature and magnetic fields
Recent results have provided strong evidence of anti-correlations between the obtained small-scale field measurements and temperature variations ($\delta T$) for a few strongly magnetic targets (Artigau et al., 2024; Cristofari et al., 2025). In this work, we investigate the impact of temperature variations on magnetic field estimates and vice versa.
To retrieve the temperature variations, we implemented a process similar to that described by Artigau et al. (2024), with the notable difference that we derived $\partial\@vec{A}/\partial T$ for each pixel from our synthetic spectra. To validate our process, we created a series of modeled observations with varying effective temperatures, following the estimate obtained with the LBL approach for EV Lac (with typical variations of less than 20 K). Our implementation allows us to successfully retrieve the temperature variations used to generate the models.
To investigate the impact of magnetic fields on temperature variation estimates, we modeled the spectra, accounting for both temperature and magnetic fields variations by taking realistic estimates for EV Lac from Cristofari et al. (2025). The average surface magnetic field strengths then ranged from 4 to 5 kG, and the effective temperatures range from 3290 to 3310 K. We found that temperature variations have a limited impact on the derived $\delta\langle B\rangle$. The inclusion of magnetic fields, however, was shown to have a much more significant impact on the derived $\delta T$ (see Fig. 2). We found that this impact is particularly significant for simulations covering the Narval/ESPaDOnS wavelength domain, where a very large number of atomic features are predicted to be sensitive to magnetic fields.
The results of this experiment show that $\delta T$ and $\delta\langle B\rangle$ estimates are not completely independent, with a greater sensitivity of $\delta T$ estimates to magnetic field variations seen for realistic conditions. We note that typical $\delta T$ are small, with a few kelvin variations leading to small changes in the observed spectra, while $\delta\langle B\rangle$ fluctuations in stars like EV Lac can lead to more visible pixel variations. Furthermore, magnetic field variations in stellar spectra lead to effects that are strongly dependent on the considered spectral line. Having both magnetically sensitive and insensitive lines in the spectral domain makes it possible to obtain robust constraints on $\delta\langle B\rangle$. A comparison between the model-driven and data-driven $\delta T$ estimates is discussed in Sect. 5.3.
@@ 4.4 Impact of atmospheric parameters on $\langle B\rangle$ variations
We investigated the impact that incorrect assumptions on the atmospheric properties of the star can have on the results. To achieve this, modeled observations were analyzed with $\@vec{F}_{k}-\@vec{F}_{0}$ computed for different effective temperatures ($T_{\rm eff}$), surface gravities ($\log{g}$), metallicities ($\rm[M/H]$), and projected rotational velocity ($v\sin{i}$).
We find that an offset in atmospheric parameters has a ‘damping’ effect on the retrieved relative field strengths, similar to the impact of short-window normalization (see Sec 4.2): the structure of the variations are typically conserved, while the amplitude of the field variations are either lower or higher than expected. Through simulations, we find that a change of $300$ K can yield a maximum difference in $\delta\langle B\rangle$ of 0.1 kG in the case of EV Lac (whose magnetic field is able to change by up to 1 kG in our simulations), comparable to the typical uncertainties reported in Cristofari et al. (2025). Similarly, changes of 0.5 dex in the assumed surface gravity or metallicity can lead to maximum differences in $\delta\langle B\rangle$ of $0.2$ kG. Changes in the assumed broadening of the spectra of up to 5 $\rm km\,s^{-1}$ typically lead to maximum changes in $\delta\langle B\rangle$ of $0.05$ kG. The typical uncertainties on $T_{\rm eff}$, $\log{g}$, $\rm[M/H]$ and $v\sin{i}$ are generally lower than 300 K, 0.5 dex or 5 $\rm km\,s^{-1}$, while the maximum differences in $\delta\langle B\rangle$ are then lower than the typical uncertainties associated with the measurements. Our process remains therefore robust to changes in atmospheric parameters and broadening, which allows us to limit the impact of incorrect stellar characterization on the results.
@@ 5.1 Application to SPIRou spectra
We applied our process to spectra recorded with SPIRou for EV Lac, DS Leo, and Barnard’s star, whose small-scale magnetic fields were previously characterized from fits of synthetic spectra to each individual observation (Cristofari et al., 2025). This allowed us to make direct comparisons between the obtained $\delta\langle B\rangle$ and the magnetic field estimates reported by Cristofari et al. (2025). For each star, we assumed atmospheric parameters that were close to those obtained by Cristofari et al. (2025), which we list in Table 1. The broadening parameters (i.e., turbulence, rotation, instrumental width) were approximated by a single Gaussian profile. While the finetuning of each of these parameters is possible, we chose to keep the assumed atmospheric parameters and broadening kernels approximate, given the relative insensitivity of the proposed method to detailed stellar characterization (see Sect. 4.4).
Table 1: Atmospheric parameters adopted in the current work. [^1] The second column shows the parameters reported by Cristofari et al. (2025). The third column presents the approximate parameters adopted for our work, corresponding the closest point in our grid of models.
@@
For all three stars, we obtained magnetic field variations that were in very good agreement with the results of Cristofari et al. (2025, as also seen in our Figs. 3 and 8), with reduced $\chi^{2}$ values between the two series of points of 1.01, 1.07, and 1.26. In addition, we obtained Pearson correlation coefficients of $0.97$, $0.89,$ and $0.58$ for EV Lac, DS Leo, and Barnard’s star, respectively. Relying on the same tools as Cristofari et al. (2025), we fit a quasi-periodic Gaussian process (QPGP) to the retrieved $\delta\langle B\rangle$ for each star. Specifically, we relied on a GP regression with a kernel, previously used in other studies (Angus et al., 2018; Fouqué et al., 2023; Cristofari et al., 2025), expressed as
$\kappa(t_{i},t_{j})=\alpha^{2}\exp\Big[-\frac{(t_{i}-t_{j})^{2}}{2l^{2}}-\frac{1}{2\beta^{2}}\sin^{2}\Big(\frac{\pi(t_{i}-t_{j})}{P_{\rm rot}}\Big)\Big]+\sigma^{2}\delta_{ij},$ (5)
where $\alpha$ is the amplitude of the GP, $\beta$ a smoothing factor, $l$ the decay time, $t_{i}-t_{j}$ the time difference between two data points, i and j, and $\sigma$ the standard deviation of an added uncorrelated white noise. Finally, $P_{\rm rot}$ is the recurrence period, which we identify as the rotation period of the star.
The posterior distributions provide constraints on the rotation period of the star that are in excellent agreement with the results of Cristofari et al. (2025) and similar long-term fluctuations were also observed. These results demonstrate our ability to estimate $\langle B\rangle$ variations that are consistent with those obtained from fit of synthetic spectra to each individual observation (see Table 7 and Figs. 12, 13, and 14). Since the proposed method to derive $\delta\langle B\rangle$ relies on solving the linear Eq. 4 and does not require us to explore the parameter space with a Markov chain Monte Carlo (MCMC), these results can be obtained faster by several orders of magnitude than the results derived by Cristofari et al. (2025).
Figure 3: Comparison between $\delta\langle B\rangle$ obtained in this work and the values of Cristofari et al. (2025). The average $\langle B\rangle$ was removed from the estimates of Cristofari et al. (2025) for comparison. The red line shows the equality. The Pearson correlation coefficient ($\rho$) is shown on the figure. Our process allows us to retrieve $\delta\langle B\rangle$ values that are in very good agreement with those obtained from fits of synthetic spectra to the data.
@@ 5.2 Application to ESPaDOnS and Narval spectra
To complete our analysis, we applied our process to optical spectra recorded with Narval and ESPaDOnS for EV Lac and DS Leo (see Sect. 2). The derived relative magnetic field strengths are well modeled by a QPGP (see Figs. 4 and 5), yielding rotation periods of $P_{\rm rot}=4.367\pm 0.023$ d and $P_{\rm rot}=14.105\pm 0.406$ d for EV Lac and DS Leo, respectively. These estimates are in very good agreement with those reported by Cristofari et al. (2025) of $P_{\rm rot}=4.362\pm 0.001$ d and $P_{\rm rot}=13.953\pm 0.088$ d, albeit with larger error bars. The larger error bars can partly be attributed to the sparsity of the available Narval and ESPaDOnS data, as a result of the observing strategies primarily focused on Zeeman Doppler Imaging reconstructions (see, e.g., Morin et al., 2008, 2011; Donati et al., 2008; Bellotti et al., 2024). This underscores the benefits of long-term spectroscopic monitoring of M dwarfs, not only for the planet search, but for the study of their magnetic fields as well. We note that DS Leo was reported to exhibit differential rotation, with distinct rotation periods at the equator and the poles ($13.37\pm 0.86$ d and $14.96\pm 1.25$ d, respectively; Hébrard et al., 2016), which could in part account for the larger error bar derived with our GP fit.
Figure 4: Best QPGP fit (green) obtained on relative small-scale magnetic field estimates derived from ESPaDOnS and Narval spectra for EV Lac. Top panel: Fit over the entire dataset. Middle and bottom: Segments of the dataset. The residuals (Res.) between data points and GP fit are shown on each panel.
Figure 5: Same as Fig. 4, but for DS Leo.